40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4416     Entry time: Fri Mar 18 17:55:58 2011
Author: Suresh 
Type: Configuration 
Category: Green Locking 
Subject: Work Plan for Y-end Aux laser installation 

A rough time-table and the various tasks are given below:

Note:  700mW NPRO sitting on AP table (Model No: 126-1064-700, Sl No. 415)  = Alberto's laser

 

 

Y-arm Aux laser installation
 1

Temperature dependence of frequency of Alberto's laser:

 a) Shifting Alberto's Laser (AL) to the PSL table and setting up a beat frequency measurement between AL and PSL

 b) Determining the frequency vs Temperature curve for the AL

Mar 21st to 25th Bryan and Suresh
2 Re-positioning the Input beam onto the IP-ANG-PD and realigning the X-arm Mar 21st to 25th Kiwamu and his 'team'  :-)
3

Repositioning the optics on the Y-end  table and relocating Alberto's laser ( at this point it will be rechiristened as Y-End-NPRO )

Mar 27th - 28th
Bryan and Suresh
4 Maximising the doubling effiiciency and obtaining the PD and QPD signals into the CDS Mar 29th - Apr 1st "
5 Aligning the Y-end green to pass through the Y-arm and locking the green to the Y arm Apr 3 - 8th "
6 Aligning the IR beam to the Y- arm and locking the Y arm to the IR Apr 10 - 15 "

 

ELOG V3.1.3-